How many pure acid must be added to a 2L of a 10% solution to obtain a mixture that is 20% acid?
Answer by addingup(3677)

Let's call the final amount x. So we want to get 0.20x
Since we have 2 at 0.10, the amount of pure acid will be x-2:
---------------------------------------------------
0.10(2)+1(x-2)= 0.20x
0.20+x-2= 0.20x
-1.8= -0.8x rearrange (for no reason, just to make it look proper)
-0.8x= -1.8 Divide both sides by -0.8 (and remember that -/-= +)
x= 2.25 liters of pure acid are need